OpenStudy (anonymous):

Which two options could be the next step in simplifying this expression? 5 • (2 + 4) A. 5 • 2 + 4 (use the distributive property) B. 5 • 2 + 5 • 4 (use the distributive property) C. (5 • 2) + 4 (use the associative property) D. 5 • 6 (use order of operations)

OpenStudy (texaschic101):

that would be : A and D distributive : 5(2 + 4) = 10 + 20 = 30 order of ops : 5(2 + 4) = 5(6) = 30